OpenStudy (anonymous):

help plz. Express each complex number in polar form. (a) 4 + 3i b) -5 – 2i

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you need two numbers to write \[4+3i\] as \[r(\cos(\theta)+i\sin(\theta)\] namely \(r\) and \(\theta\)

OpenStudy (misty1212):

\(r\) is easiest, it is the absolute value or \(\sqrt{a^2+b^2}\)

OpenStudy (misty1212):

then for \(\theta\) depending on what quadrant you are in, it may be \[\theta=\tan^{-1}(\frac{b}{a})\]

OpenStudy (misty1212):

that will work for \(4+3i\) it will not work for \(-5-2i\)because you are in quadrant 3 there